# Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

### v1

@@ 6 Results
Table 2: Comparative Analysis of FDS Component Effects Across PACS Dataset Domains
Module Accuracy (%) Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 + Image Generation (No Interpol.) 89.87 $\pm$0.1 85.59 $\pm$0.6 97.50 $\pm$0.3 83.07 $\pm$0.4 89.01 $\pm$0.4 + Image Interpolation 91.38 $\pm$0.2 85.20 $\pm$0.6 97.73 $\pm$0.1 84.27 $\pm$0.9 89.65 $\pm$0.4 + Filtering Mechanism 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
We first compare the accuracy of our approach against state-of-art DG methods, across four benchmarks. We then present a detailed analysis investigating several key aspects of our approach, including the effectiveness of each proposed component, its regularization capabilities, and a comparison of different mixing strategies.
6.1 Comparison with the state of the art
In Table 1, we compare our approach with recent methods for domain generalization as outlined in the DomainBed framework [51]. Our method, when added to the standard ERM classifier, shows an improved accuracy of 3.1% on the PACS dataset using the ResNet-50 model. On the VLCS dataset, our method sees a 3.23% increase in accuracy over the standard ERM. For the OfficeHome dataset, our method outperforms the ERM baseline by 4.57%. To test the strength of our method, we also applied it to SWAD, which is a strong baseline for domain generalization. Here, our method improves the performance on the PACS dataset by 2.39%, on the VLCS dataset by 0.95%, and on the OfficeHome dataset by 3.17%. Detailed results of accuracy on each domain of all datasets can be found in Supplemental Material.
@@ 6.2 Further Analysis
Ablation Study on different Components.  To evaluate the impact of each component on out-of-distribution (OOD) generalization, we systematically introduce each module and observe the enhancements. As detailed in Table 2 first incorporating domain-specific synthetic samples, generated without interpolation, yields a $0.91\%$ gain over the baseline. This increment validates our diffusion model’s proficiency in capturing and replicating the class-specific distributions within domains, thus refining OOD performance through enhanced density estimation near original samples. Subsequently, integrating images from pseudo-novel domains via our interpolation strategy leads to a further $0.5\%$ enhancement, underscoring the mechanism’s effectiveness in connecting distinct domains. Lastly, applying our entropy-based filtering to eliminate overly simplistic images results in a significant $2.4\%$ improvement against the SWAD model, a robust benchmark. These outcomes collectively underscore the efficacy of our approach in fortifying OOD generalization.
@@
Scaling Law of Data Size.  Next, we assess the impact of the selected data size, denoted as $N_{L}$, on the augmented dataset’s efficacy. We consider the PACS dataset for this analysis, where the average number of images per class is 570. We systematically explore varying scales relative to this average class size, aiming to discern the optimal dataset size for enhancing out-of-domain (OOD) generalization. Our findings, detailed in Figure 4, reveal an initial improvement in OOD generalization with increased data size. However, excessively enlarging the dataset size begins to diminish the benefits of our filtering mechanism, as it incorporates a broader array of samples, including those that are overly simplistic and not conducive to model improvement.
@@
Table 3: Impact of Filtering Strategy Components on PACS Dataset Accuracy
Method Accuracy (%) PACS VLCS OfficeHome Avg. Baseline (SWAD [24]) 88.11 $\pm$0.4 78.87 $\pm$0.5 70.34 $\pm$0.4 79.11 $\pm$0.4 Filtered Based on Entropy 90.50 $\pm$0.4 79.33 $\pm$0.8 71.97 $\pm$0.3 80.60 $\pm$0.5 + Reject Extreme Transform. 90.50 $\pm$0.3 79.73 $\pm$0.5 73.51 $\pm$0.5 81.25 $\pm$0.4
Figure 3: Accuracy Comparison on the PACS Dataset: Random Selection vs. Proposed Filtering Strategy of FDS
Figure 4: Impact of varying scales of sample size $N_{L}$ relative to the average number of images per class on Accuracy in the PACS Dataset Analysis.
Mixing Strategies.  Our ablation study, presented in Table 4, reveals that both Noise Level Interpolation and Condition Level Interpolation significantly outperform the baseline, yet their effectiveness varies with the dataset’s attributes and the nature of the domain shift. Noise Level Interpolation is optimal for minimal domain shifts, focusing on adjusting the noise aspect to bridge domain gaps. However, it falls short in scenarios with substantial domain differences, such as the transition to Cartoon or Sketch, where the source and target domains diverge significantly. In these cases, Condition Level Interpolation proves more advantageous, offering a robust mechanism for navigating complex domain shifts by manipulating higher-level semantic representations. This strategy is particularly effective when conventional noise adjustments are insufficient for achieving meaningful domain traversal.
Table 4: Out-of-Distribution Generalization Performance Using Various Interpolation Strategies of FDS
Strategy Accuracy (%) Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 Noise Level Interpol. 91.95 $\pm$0.3 83.23 $\pm$0.1 97.56 $\pm$0.1 85.85 $\pm$0.3 89.65 $\pm$0.2 Condition Level Interpol. 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
6.2.1 In-Domain Regularization Effect.
In this section, we study the impact of incorporating synthetic images produced by our method on the accuracy of the network when tested on the same domains on which it was trained. As indicated in Table 5, FDS approach surpasses the baseline accuracy within this context. This affirms the value of our method, not only in out-of-distribution (OOD) conditions, but also in standard in-domain validation, aligning with the principles of Vicinal Risk Minimization (VRM) [57]. Hence, our method may also be viewed as a regularization strategy, suitable for a broad spectrum of applications.
Table 5: Effect of FDS on In-Domain Accuracy: PACS Dataset Analysis
@@
6.3 Stability Analysis
In this section, we demonstrate the performance of our model across different stages of training within two domains of the PACS dataset, depicted in Figure 5. It is important to note that these test accuracies were not used in the selection of the best-performing model mentioned in earlier sections and all of our experiments follow leave-one-out settings suggested by DomainBed protocol. A notable issue with domain shift, as evidenced by our findings, is the substantial and irregular fluctuations in test accuracy during training. This typically necessitates a comprehensive exploration of hyperparameter settings. For context, as reported in Table 6, training an ERM model within the DomainBed framework required 4,142 experiments to determine the most effective hyperparameters. In contrast, our approach yielded more consistent test accuracies over time, removing the need for such exhaustive hyperparameter searches. Consequently, we were able to achieve and report our results using a single, default set of hyperparameters. Details of hyperparameter search space for each framework is included in supplemental materials.
Figure 5: Test Accuracy Across Training Intervals: Comparison between ERM (Top Row) vs. FDS (Bottom Row) in Art and Sketch Domains of PACS
Framework #Experiments DomainBed [51] 4142 SWAD [24] 396 Ours 1
Table 6: Size of HP search space in different frameworks
@@ 7 Conclusion
This work presented FDS, a domain generalization (DG) technique that leverages diffusion models for domain mixing, generating a diverse set of images to bridge the domain gap between source domains distribution. We also proposed an entropy-based filtering strategy enriches the pseudo-novel generated set with images that test the limits of classifiers trained on original data, thereby boosting generalization. Our extensive experiments across multiple benchmarks demonstrate that our method not only surpasses existing DG techniques but also sets new records for accuracy. Our analysis indicate that our approach contributes to more stable training processes when confronted with domain shifts and serves effectively as a regularization method in in-domain contexts. Notably, our technique consistently enhances performance across diverse scenarios, from realistic photos to sketches, without necessitating extensive hyperparameter tuning commonly associated with other domain generalization strategies. Moreover, while we exploited our trained diffusion model for covering the domain gap, more sophisticated techniques could be considered. For instance, future work could investigate the idea of generating images by extrapolate in domain space to explore pseudo-novel distributions.
@@ Appendix B Additional Ablation
Table 9: Analysis of the impact of utilizing different numbers/combinations of generated domains on final model performance across the PACS dataset domains. For definitions of each augmented domain (ID0, ID1, ID2), see Table 10.
@@
Appendix C Detailed Results
Here are the comprehensive tables containing all the detailed information that was summarized in the main paper.
Table 11: Deatailed Comparison of FDS with the state-of-the-art on PACS dataset. numbers marked are reproduced in-house.
@@ Appendix C Detailed Results
Method Source Domains Art Cartoon Photo Sketch Avg. CDANN 84.6 $\pm$1.8 75.5 $\pm$0.9 96.8 $\pm$0.3 73.5 $\pm$0.6 82.60 $\pm$0.9 MASF 82.9 80.5 95.0 72.3 82.7 DMG 82.6 78.1 94.5 78.3 83.4 IRM 84.8 $\pm$1.3 76.4 $\pm$1.1 96.7 $\pm$0.6 76.1 $\pm$1.0 83.50 $\pm$1.0 MetaReg 87.2 79.2 97.6 70.3 83.6 DANN 86.4 $\pm$0.8 77.4 $\pm$0.8 97.3 $\pm$0.4 73.5 $\pm$2.3 83.65 $\pm$1.1 ERM 85.7 $\pm$0.6 77.1 $\pm$0.8 97.4 $\pm$0.4 76.6 $\pm$0.7 84.20 $\pm$0.6 GroupDRO 83.5 $\pm$0.9 79.1 $\pm$0.6 96.7 $\pm$0.3 78.3 $\pm$2.0 84.40 $\pm$0.9 MTL 87.5 $\pm$0.8 77.1 $\pm$0.5 96.4 $\pm$0.8 77.3 $\pm$1.8 84.58 $\pm$0.9 I-Mixup 86.1 $\pm$0.5 78.9 $\pm$0.8 97.6 $\pm$0.1 75.8 $\pm$1.8 84.60 $\pm$0.8 MMD 86.1 $\pm$1.4 79.4 $\pm$0.9 96.6 $\pm$0.2 76.5 $\pm$0.5 84.65 $\pm$0.7 VREx 86.0 $\pm$1.6 79.1 $\pm$0.6 96.9 $\pm$0.5 77.7 $\pm$1.7 84.93 $\pm$1.1 MLDG 85.5 $\pm$1.4 80.1 $\pm$1.7 97.4 $\pm$0.3 76.6 $\pm$1.1 84.90 $\pm$1.1 ARM 86.8 $\pm$0.6 76.8 $\pm$0.5 97.4 $\pm$0.3 79.3 $\pm$1.2 85.08 $\pm$0.6 RSC 85.4 $\pm$0.8 79.7 $\pm$1.8 97.6 $\pm$0.3 78.2 $\pm$1.2 85.23 $\pm$1.0 Mixstyle 86.8 $\pm$0.5 79.0 $\pm$1.4 96.6 $\pm$0.1 78.5 $\pm$2.3 85.23 $\pm$1.1 ER 87.5 79.3 98.3 76.3 85.3 pAdaIN 85.8 81.1 97.2 77.4 85.4 EISNet 86.6 81.5 97.1 78.1 85.8 CORAL 88.3 $\pm$0.2 80.0 $\pm$0.5 97.5 $\pm$0.3 78.8 $\pm$1.3 86.15 $\pm$0.6 SagNet 87.4 $\pm$1.0 80.7 $\pm$0.6 97.1 $\pm$0.1 80.0 $\pm$0.4 86.30 $\pm$0.5 DSON 87.0 80.6 96.0 82.9 86.6 SWAD 89.3 $\pm$0.2 83.4 $\pm$0.6 97.3 $\pm$0.3 82.5 $\pm$0.5 88.13 $\pm$0.4 ERM 86.94 $\pm$0.6 80.21 $\pm$0.7 96.61 $\pm$0.4 74.45 $\pm$2.9 84.30 $\pm$1.1 ERM + FDS (ours) 86.27 $\pm$3.1 84.24 $\pm$0.7 97.08 $\pm$0.9 81.91 $\pm$1.9 87.38 $\pm$1.6 SWAD 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$0.1 88.11 $\pm$0.4 SWAD + FDS (ours) 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
Table 12: Deatailed Comparison of FDS with the state-of-the-art on VLCS dataset. numbers marked are reproduced in-house.

### v2

@@ 6 Results
We first compare the performance of our FDS approach against SOTA DG methods across three benchmarks. We then present a detailed analysis investigating several key aspects of our approach, including the effectiveness of each proposed component, a comparison of different mixing strategies, its regularization capabilities, domain diversity visualization and quantification, the impact of data size, the efficacy of our filtering mechanism, and stability analysis during training. Additional analysis, visualizations, and detailed tables can be found in the Supplementary Material.
6.1 Comparison with the State-of-the-art
In Table 1, we compare our approach with recent methods for domain generalization as outlined in the DomainBed framework [44]. Our method, when added to the baseline ERM classifier, shows an impressive improved accuracy of 4.5% on the PACS dataset using the ResNet-50 model. On the VLCS dataset, our method sees a 3.6% increase in accuracy over the ERM. For the OfficeHome dataset, our method outperforms the ERM baseline by 6.5%. To test the strength of our method, we also applied it to SWAD, which is a the baseline for weight averaging (WA) domain generalization methods. Here, our method improves the performance by 2.4%, 0.8%, and 3.2% on the PACS, VLCS, and OfficeHome datasets, respectively.
Furthermore, when comparing with previous SOTA methods, our method outperforms CDGA [55], by 0.3% on PACS, 0.2% on VLCS, and 2.9% on OfficeHome. In the context of weight averaging methods, our approach surpasses DIWA [68], which trains multiple independent models, by 1.7% on PACS, 0.6% on VLCS, and 2.5% on OfficeHome, setting a new benchmark for domain generalization. Please refer to the Supplementary Material for the full, detailed results for each dataset and its domains.
@@ 6.2 Further Analysis
For this section, we use the SWAD baseline to analyze the performance of our proposed method FDS, due to its stable performance as a WA method. Final settings are applied to the ERM baseline. All analyses in this section use the PACS dataset and the SWAD baseline unless otherwise stated.
Table 2: Comparative analysis of FDS component effects on accuracy (%) across PACS dataset domains. "Basic" refers to generation without interpolation or filtering.
Module Target Domains Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 + Basic 89.87 $\pm$0.1 85.59 $\pm$0.6 97.50 $\pm$0.3 83.07 $\pm$0.4 89.01 $\pm$0.4 + Interpolation 91.38 $\pm$0.2 85.20 $\pm$0.6 97.73 $\pm$0.1 84.27 $\pm$0.9 89.65 $\pm$0.4 + Filtering 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
@@
Ablation Study on Different Components.  To evaluate the impact of each component on generalization, we systematically introduce each module and observe the enhancements. As detailed in Table 2, first incorporating domain-specific synthetic samples, generated without interpolation or filtering (Basic), yields a $0.9\%$ gain over the baseline. This increment validates our diffusion model’s proficiency in capturing and replicating the class-specific distributions within domains, thus refining OOD performance through enhanced density estimation near original samples. Subsequently, integrating images from pseudo-novel domains via our interpolation strategy leads to a further $0.5\%$ enhancement, underscoring the mechanism’s effectiveness in connecting distinct domains. Lastly, applying our entropy-based filtering to eliminate overly simplistic images and images with diminished semantic relevance results in a significant $2.4\%$ improvement against the SWAD model, a robust benchmark. These outcomes collectively underscore the efficacy of our approach in improving OOD generalization.
Figure 3: Impact of varying scales of sample size $N_{L}$ relative to the average number of images per class on PACS dataset.
Figure 4: Impact of using Random Selection vs. Proposed Filtering Strategy of FDS on PACS accuracy (%).
Table 3: Impact of different interpolation strategies of FDS on PACS accuracy (%).
Strategy Target Domains Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 Noise Level Interpol. 91.95 $\pm$0.3 83.23 $\pm$0.1 97.56 $\pm$0.1 85.85 $\pm$0.3 89.65 $\pm$0.2 Condition Level Interpol. 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3 Both 91.13 $\pm$0.4 82.98 $\pm$0.2 97.90 $\pm$0.2 85.62 $\pm$0.8 89.41 $\pm$0.4
Mixing Strategies.  Our ablation study, presented in Table 3, reveals that both Noise Level Interpolation and Condition Level Interpolation significantly outperform the baseline, yet their effectiveness varies with the dataset’s attributes and the nature of the domain shift. Noise Level Interpolation is optimal for minimal domain shifts, focusing on adjusting the noise aspect to bridge domain gaps. However, it falls short in scenarios with substantial domain differences, such as the transition to Cartoon or Sketch, where the source and target domains diverge significantly. In these cases, Condition Level Interpolation proves more advantageous, offering a robust mechanism for navigating complex domain shifts by manipulating higher-level semantic representations. When applying both methods simultaneously, the performance did not improve compared to using Condition Level Interpolation alone, possibly due to the increased complexity. Nonetheless, it still performed better than the baseline. Based on these results, we use Condition Level Interpolation as our final interpolation method.
Figure 5: t-SNE plots showcasing the original and generated samples for the "Art", "Photo", and "Sketch" source domains and the "giraffe" class in the PACS dataset using our method. Features were extracted using the CLIP vision encoder [71].
@@
Impact of Sample Size.  Next, we assess the impact of the selected data size, denoted as $N_{L}$, on final performance. We consider the PACS dataset for this analysis, where the average number of images per class is 570. We systematically explore varying scales relative to this average class size, aiming to discern the optimal dataset size for enhancing OOD generalization. Our findings, detailed in Figure 4, reveal an initial improvement in OOD generalization with increased data size. However, excessively enlarging the dataset size begins to diminish the benefits of our filtering mechanism, as it incorporates a broader array of samples, including those that are overly simplistic and not conducive to model improvement.
@@
Table 4: Impact of filtering strategy components of FDS on accuracy (%), using three benchmarks.
Method PACS VLCS Office  Avg. Baseline (SWAD [24]) 88.11 $\pm$0.4 78.87 $\pm$0.5 70.34 $\pm$0.4 79.11 Filtered Based on Entropy 90.50 $\pm$0.4 79.33 $\pm$0.8 71.97 $\pm$0.3 80.60 + Reject Semantic Loss 90.50 $\pm$0.3 79.73 $\pm$0.5 73.51 $\pm$0.5 81.25
Table 5: Domain diversity metric [72] between source domains and the target domain of the PACS dataset. "Basic" refers to generation without interpolation or filtering.
Data Target Domains Art Cartoon Photo Sketch Avg. Original PACS 0.39 0.87 0.54 0.99 0.70 Basic 0.54 0.72 0.45 1.01 0.68 FDS 0.44 0.58 0.49 0.92 0.61
Domain Diversity Visualization.  To illustrate the effectiveness of our method, we present t-SNE plots of original PACS dataset samples and those generated by FDS in Figure 5. The t-SNE plots show distinct clusters for original source domains (here Art, Photo, Sketch) and highlight how the generated samples clearly bridge the gaps between these clusters, thus enhancing domain diversity. This expanded diversity is crucial for improving the generalization capabilities of models, as it ensures a broader spectrum of data distributions in the training set. By providing a continuous representation of the domain space, our method facilitates smoother transitions and better prepares models to handle unseen domains, ultimately contributing to more robust performance in real-world applications.
Domain Diversity Quantification.  We further validate the effectiveness of our method using a domain diversity metric based on the methodology proposed in [72]. This metric quantifies the diversity shift between the source domains and the target domain, providing valuable insight into how the newly generated domains using our method can improve generalization on unseen domains. Table 5 presents the domain diversity metric for the original PACS dataset, comparing it with samples generated using the diffusion model in its basic form (no interpolation and no filter) and our FDS method (including interpolation and filtering). This table demonstrates that our FDS method results in a lower diversity metric compared to using the original PACS dataset or the samples generated with basic generation, suggesting that our method more effectively promotes generalization to new, unseen domains.
Table 6: Impact of FDS on in-domain PACS accuracy (%). ’A’, ’C’, ’P’, and ’S’ show ’Art’, ’Cartoon’, ’Photo’, and ’Sketch’ domains.
@@
In-Domain Regularization Effect.  In this section, we study the impact of incorporating generated images produced by our method on the accuracy of the network when tested on the same domains on which it was trained. As indicated in Table 6, our FDS approach surpasses the baseline accuracy within this context. This affirms the value of our method, not only in OOD conditions, but also in standard in-domain validation, aligning with the principles of Vicinal Risk Minimization (VRM) [57]. Hence, our method may also be viewed as a regularization strategy, suitable for a broad spectrum of applications.
Figure 6: Accuracy (%) across training steps: Comparison between ERM (top row) vs. FDS (bottom row) in "Art" and "Sketch" domains of PACS dataset.
Stability Analysis.  In this section, we demonstrate the performance of our model across different stages of training within two domains of the PACS dataset, depicted in Figure 6. It is important to note that these test accuracies were not used in the selection of the best-performing model mentioned in earlier sections and all of our experiments follow leave-one-out settings suggested by DomainBed. The results indicate that our model achieves higher stability and better mean accuracy with lower standard deviation compared to the ERM trained on original data. Note that we cannot plot the figures for SWAD since it is a WA of ERM and does not have individual training curves. These results demonstrate the robustness and stability of our model during training, which is crucial for domain generalization algorithms.
@@ 7 Conclusion
This work presented FDS, a domain generalization (DG) technique that leverages diffusion models for domain mixing, generating a diverse set of images to bridge the domain gap between source domains distribution. We also proposed an entropy-based filtering strategy enriches the pseudo-novel generated set with images that test the limits of classifiers trained on original data, thereby boosting generalization. Our extensive experiments across multiple benchmarks demonstrate that our method not only surpasses existing DG techniques but also sets new records for accuracy. Our analysis indicate that our approach contributes to more stable training processes when confronted with domain shifts and serves effectively as a regularization method in in-domain contexts. Notably, our technique consistently enhances performance across diverse scenarios, from realistic photos to sketches. Moreover, while we exploited our trained diffusion model for covering the domain gap, more sophisticated techniques could be considered. For instance, future work could investigate the idea of generating images by extrapolate in domain space to explore pseudo-novel distributions.
FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization - Supplementary Material
@@ Appendix B Additional Ablation
Table 9: Analysis of the impact of utilizing different numbers/combinations of generated domains on final model performance across the PACS dataset domains (Leave-one-out accuracy). For definitions of each augmented domain (ID0, ID1, ID2), see Table 10.
@@
Appendix C Oracle Results
In addition to leave-one-out setting, where the validation set is selected from the training domains, some studies also report the results of oracle (test-domain validation set). This can be particularly useful for understanding the potential of a method when domain knowledge is available. In this section, we compare our method (FDS+ERM) with the state-of-the-art results, as shown in Table 11. It is important to note that no Weight Averaging (WA) methods reported their oracle results within the DomainBed framework for a fair comparison. Therefore, we only train and report our ERM results here. Our proposed method, FDS+ERM, demonstrates superior performance across multiple benchmarks. Specifically, it achieves an average accuracy of 81.2%, outperforming all other methods. On the PACS dataset, FDS+ERM attains the highest accuracy of 89.7%, with significant improvements in the VLCS and OfficeHome datasets as well, achieving accuracies of 82.0% and 71.8% respectively. In addition to leave-one-out setting, these results also highlight the effectiveness of our approach in enhancing the performance under the oracle setting.
Table 11: Oracle (test-domain validation set) accuracy (%) results on the PACS, VLCS, and OfficeHome benchmarks. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
Method Aug. PACS VLCS OfficeHome  Avg. Standard Methods ERM (baseline[44] 86.7 $\pm$0.3 77.6 $\pm$0.3 66.4 $\pm$0.5 76.9 ERM (reproduced) 86.6 $\pm$0.8 79.8 $\pm$0.4 68.4 $\pm$0.3 78.3 IRM [60] 84.5 $\pm$1.1 76.9 $\pm$0.6 63.0 $\pm$2.7 74.8 GroupDRO [47] 87.1 $\pm$0.1 77.4 $\pm$0.5 66.2 $\pm$0.6 76.9 Mixup [61] 86.8 $\pm$0.3 78.1 $\pm$0.3 68.0 $\pm$0.2 77.6 CORAL [62] 87.1 $\pm$0.5 77.7 $\pm$0.2 68.4 $\pm$0.2 77.7 MMD [63] 87.2 $\pm$0.1 77.9 $\pm$0.1 66.2 $\pm$0.3 77.1 DANN [64] 85.2 $\pm$0.2 79.7 $\pm$0.5 65.3 $\pm$0.8 76.7 SagNet [65] 86.4 $\pm$0.4 77.6 $\pm$0.1 67.5 $\pm$0.2 77.2 RSC [23] 86.2 $\pm$0.5 66.5 $\pm$0.6 SelfReg [41] 86.7 $\pm$0.8 78.2 $\pm$0.1 68.1 $\pm$0.3 77.7 Fishr [73] 85.8 $\pm$0.6 78.2 $\pm$0.2 66.0 $\pm$2.9 76.7 CDGA [55] 89.6 $\pm$0.3 80.9 $\pm$0.1 68.8 $\pm$0.3 79.3 ERM + FDS (ours) 89.7 $\pm$0.8 82.0 $\pm$0.1 71.8 $\pm$0.9 81.2
Appendix D Detailed Results
Here we present the comprehensive tables containing all the detailed information that was summarized in the main paper. The leave-one-out performance (train-domain validation set) across different domains of PACS, VLCS, and OfficeHome datasets are detailed in the tables 12, 13, and 14, respectively. Additionally, the oracle (test-domain validation set) accuracy results for the PACS, VLCS, and OfficeHome benchmarks are detailed in Table 15, 16, and 17, respectively.
Table 12: Leave-one-out accuracy (%) results on the PACS dataset. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
@@ Appendix D Detailed Results
Method Aug. Target Domains Art Cartoon Photo Sketch Avg. Standard Methods ERM (baseline) [44] 84.7 $\pm$0.4 80.8 $\pm$0.6 97.2 $\pm$0.3 79.3 $\pm$1.0 85.5 $\pm$0.2 ERM (reproduced) 86.9 $\pm$0.6 80.2 $\pm$0.7 96.6 $\pm$0.4 74.5 $\pm$2.9 84.3 $\pm$1.1 IRM [60] 84.8 $\pm$1.3 76.4 $\pm$1.1 96.7 $\pm$0.6 76.1 $\pm$1.0 83.5 $\pm$0.8 GroupDRO [47] 83.5 $\pm$0.9 79.1 $\pm$0.6 96.7 $\pm$0.3 78.3 $\pm$2.0 84.4 $\pm$0.8 Mixup [61] 86.1 $\pm$0.5 78.9 $\pm$0.8 97.6 $\pm$0.1 75.8 $\pm$1.8 84.6 $\pm$0.6 CORAL [62] 88.3 $\pm$0.2 80.0 $\pm$0.5 97.5 $\pm$0.3 78.8 $\pm$1.3 86.2 $\pm$0.3 MMD [63] 86.1 $\pm$1.4 79.4 $\pm$0.9 96.6 $\pm$0.2 76.5 $\pm$0.5 84.6 $\pm$0.5 DANN [64] 86.4 $\pm$0.8 77.4 $\pm$0.8 97.3 $\pm$0.4 73.5 $\pm$2.3 83.6 $\pm$0.4 MLDG [15] 85.5 $\pm$1.4 80.1 $\pm$1.7 97.4 $\pm$0.3 76.6 $\pm$1.1 84.9 $\pm$1.1 VREx [74] 86.0 $\pm$1.6 79.1 $\pm$0.6 96.9 $\pm$0.5 77.7 $\pm$1.7 84.9 $\pm$1.1 ARM [75] 86.8 $\pm$0.6 76.8 $\pm$0.5 97.4 $\pm$0.3 79.3 $\pm$1.2 85.1 $\pm$0.6 SagNet [65] 87.4 $\pm$1.0 80.7 $\pm$0.6 97.1 $\pm$0.1 80.0 $\pm$0.4 86.3 $\pm$0.2 RSC [23] 85.4 $\pm$0.8 79.7 $\pm$1.8 97.6 $\pm$0.3 78.2 $\pm$1.2 85.2 $\pm$0.9 Mixstyle [20] 86.8 $\pm$0.5 79.0 $\pm$1.4 96.6 $\pm$0.1 78.5 $\pm$2.3 85.2 $\pm$0.3 mDSDI [66] 87.7 $\pm$0.4 80.4 $\pm$0.7 98.1 $\pm$0.3 78.4 $\pm$1.2 86.2 $\pm$0.2 SelfReg [41] 87.9 $\pm$1.0 79.4 $\pm$1.4 96.8 $\pm$0.7 78.3 $\pm$1.2 85.6 $\pm$0.4 Fishr [73] 88.4 $\pm$0.2 78.7 $\pm$0.7 97.0 $\pm$0.1 77.8 $\pm$2.0 85.5 $\pm$0.5 DCAug [59] 88.5 $\pm$0.8 78.8 $\pm$1.5 96.3 $\pm$0.1 80.8 $\pm$0.5 86.1 $\pm$0.7 CDGA [55] 89.1 $\pm$1.0 82.5 $\pm$0.5 97.4 $\pm$0.2 84.8 $\pm$0.9 88.5 $\pm$0.5 ERM + FDS (ours) 90.7 $\pm$0.9 84.2 $\pm$0.6 97.2 $\pm$0.1 83.0 $\pm$0.4 88.8 $\pm$0.1 WA Methods SWAD (baseline) [24] 89.3 $\pm$0.2 83.4 $\pm$0.6 97.3 $\pm$0.3 82.5 $\pm$0.5 88.1 $\pm$0.1 SWAD (reproduced) 89.5 $\pm$0.2 83.7 $\pm$0.4 97.3 $\pm$0.2 82.1 $\pm$0.1 88.1 $\pm$0.4 SelfReg SWA [41] 85.9 $\pm$0.6 81.9 $\pm$0.4 96.8 $\pm$0.1 81.4 $\pm$0.6 86.5 $\pm$0.3 DNA [67] 89.8 $\pm$0.2 83.4 $\pm$0.4 97.7 $\pm$0.1 82.6 $\pm$0.2 88.4 $\pm$0.1 DiWA [68] 90.1 $\pm$0.6 83.3 $\pm$0.6 98.2 $\pm$0.1 83.4 $\pm$0.4 88.8 $\pm$0.4 TeachDCAug [59] 89.6 $\pm$0.0 81.8 $\pm$0.5 97.7 $\pm$0.0 84.5 $\pm$0.2 88.4 $\pm$0.2 SWAD + FDS (ours) 91.8 $\pm$0.3 86.0 $\pm$0.8 98.1 $\pm$0.2 86.1 $\pm$0.1 90.5 $\pm$0.3
Table 13: Leave-one-out accuracy (%) results on the VLCS dataset. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.